What is the value of y(π

2


)

y(π2) where y

′′

−2y

+y=xe

x

sinx

y″−2y′+y=xexsin⁡x ; y(0)=0

y(0)=0 and y

(0)=1

y′(0)=1 ?


1
Expert's answer
2021-09-27T15:41:13-0400

"k^2-2k+1=0"

"k_{1,2}=1"

"Y=c_1e^x+c_2xe^x"


"y(x)=c_1(x)e^x+c_2(x)xe^x"


"y_1(x)\frac{dc_1(x)}{dx}+y_2(x)\frac{dc_2(x)}{dx}=0"


"\frac{dy_1(x)}{dx}\frac{dc_1(x)}{dx}+\frac{dy_2(x)}{dx}\frac{dc_2(x)}{dx}=xe^xsinx"


"e^x\frac{dc_1(x)}{dx}+xe^x\frac{dc_2(x)}{dx}=0"


"\frac{d}{dx}e^x\frac{dc_1(x)}{dx}+\frac{d}{dx}(xe^x)\frac{dc_2(x)}{dx}=xe^xsinx"


"\frac{d}{dx}e^x\frac{dc_1(x)}{dx}+\frac{d}{dx}(xe^x+e^x)\frac{dc_2(x)}{dx}=xe^xsinx"



"\frac{d}{dx}c_1(x)=-x^2sinx"


"\frac{d}{dx}c_2(x)=xsinx"


"c_1(x)=c_3+\int(-x^2sinx)dx"


"c_2(x)=c_4+\int xsinxdx"


"c_1(x)=c_3+x^2cosx-2xsinx-2cosx"

"c_2(x)=c_4-xcosx+sinx"


"y(x)=c_3e^x+c_4xe^x-xe^xsinx-2e^xcosx"


"y(0)=c_3-2=0\implies c_3=2"


"y'(x)=c_3e^x+c_4xe^x+c_4e^x-e^x(sinx+xcosx)-xe^xsinx-2e^xcosx+2e^xsinx"

"y'(0)=c_3+c_4-2=1\implies c_4=1"


"y(x)=2e^x+xe^x-xe^xsinx-2e^xcosx"


"y(\pi/2)=9.62"